# Unveiling Topological Structures from Language: A Comprehensive Survey of Topological Data Analysis Applications in NLP

### v3

@@ 3.1.1 Semantic Topological Space
A semantic topological space is a conceptual framework used to represent and analyze the relationships between meanings (semantics) of words, phrases, or other linguistic units in a topological or shape structure. This representation involves mapping these units into a mathematical space where the distance or structure between them reflects semantic similarity or other relationships (i.e., Euclidean space $\to$ Topological space).
Due to the limitations of traditional clustering algorithms in capturing semantic relationships within texts, Chiang (2007) utilizes TDA and represents semantic space as a simplicial complex of Euclidean space. This technique captures the coherent topics of a document in the connected component dimension ($\beta_{0}$), representing the semantic topological space Chiang (2007). Similarly, Karlgren et al. (2014) visualizes the topological semantic space of text using Mapper which identifies the topical density of the space. To capture topological properties, they train two semantic spaces in a specific topical domain Karlgren et al. (2014). One space was trained only on articles of similar topics, and the other on introductory paragraphs of those same articles. Findings reveal that clusters of main concepts remained close for the space trained only on articles of similar topics. For the other topological space, the main concepts were randomly distributed Karlgren et al. (2014).
In addition, Cavaliere et al. (2017) extracts main concepts from the texts by probing the context-aware semantic topological space built with simplicial complexes. Finally, Wagner et al. (2012) uses TF-IDF to numerically represent the top 10-50 words in a corpus and build a topological space that analyzes the structure of similarities within several documents. This topological space is built using discrete Morse theory and persistent homology to find meaningful topological patterns Wagner et al. (2012).
@@ 3.1.2 Syntactic Topological Space
Therefore, Port et al. (2018) analyzes how syntactic parameters are distributed over different language families, including Indo-European, Niger-Congo, Austronesian, and Afro-Asiatic families. , and features in $\beta_{1}$ capture syntactic differences between branches of families of languages, as well as the syntactic influences between them Port et al. (2018). Port et al. (2018) show that the three persistent connected components ($\beta_{0}$) in the Niger-Congo family represents its three subfamilies - Mande, Atlantic-Congo, and Kordofania. The syntactic topological structures of these languages also reveal the historical linguistic phenomena that the Hellenic branch played a role in the historical development of the Indo-European languages Port et al. (2018).
Similarly, Port et al. (2022) probes the interpretability of the syntactic topological space. This is done by incorporating more explanation of linguistic structures by introducing so-called homoplasy phenomena to explain persistent loops further. Homoplasy phenomena in syntax are observed when dissimilar languages exhibit syntactic similarities Port et al. (2022). Findings reveal that the Indo-European family languages - Czech, Lithuanian, Middle Dutch, and Swiss German have the same homoplasy phenomena Port et al. (2022).
Insight (Syntactic Space). The syntactic topological space captures the syntactic structure of language (i.e., grammar, etc.) from a topological lens. Using this framework, researchers confirm linguistic phenomena in language families and subfamilies by exploring the syntactic relationship between languages. Thus, such a framework can be further utilized to discover novel linguistic phenomena.
@@ 3.2 Topology of Topic Evolution
Thus, Sami and Farrahi (2017) utilizes TDA to visualize the relationship between words in a text block, words in a corpus, and text blocks in a corpus. They visualize both local context (i.e., each text block in a set of sentences), and global context (i.e., occurrence of extracted words in the corpus) features. These features are extracted by using the circular topology to represent words. Finally, findings reveal that using the circular topology in 2D space, core words from the corpus stay close to the center, and the explanatory words remain close to the circle’s periphery.
@@ 3.3 Topological “Shape” of Words
Draganov and Skiena (2024) captures the “shape” of words for several languages by comparing the phylogenies or evolutionary history of language in the Indo-European language family. Initially, numerically representing the texts with FastText, they use persistent homology to construct language phylogenetic trees for over 81 Indo-European languages. Experiments reveal that: (1) the shape of the word embedding of a language carries historical and structural information, similar to Port et al. (2018, 2022)’s findings; and (2) TDA methods can successfully capture aspects of the shape of language embeddings Draganov and Skiena (2024). Next, Fitz (2022) introduces a novel terminology - word manifold, which is a simplicial complex, whose topological space captures grammatical structure expressed by the corpus. This is done by implementing a technique for generating topological structure directly from strings of words. Experiments reveal that the homotopy type of the word manifold is also influenced by linguistic structure Fitz (2022). Similarly, Fitz et al. (2024) measures the topological complexity (known as perforation) of hidden representation of Transformer-based models to understand their topological shapes.
Similar to Port et al. (2018, 2022), Dong (2024) extracts the topological shapes of languages. Specifically South American languages - the Nuclear-Macro-Jê (NMJ) and Quechuan families using TDA. By using techniques like multiple correspondence analysis (MCA) for dimension reduction of the categorical-valued dataset and persistent homology, Dong (2024) visualizes each language in the selected families as a point cloud. This forms the topological shape of the South American languages, such that languages close together are more similar. By comparing the topological shapes of the languages, it is observed that there are major distinctions between the Jê-proper and the non-Jê-proper languages, as well as the northern and southern Quechuan languages Dong (2024). Finally, Bouazzaoui et al. (2021) explores the topological similarity of the shapes of two writing systems - Tifinagh and Phoenician scripts.
@@
Figure 4: Illustration of the TDA feature extraction pipeline in NLP with three steps: 1-extracting numerical representations, 2-reformatting for TDA’s inputs, and 3-extracting TDA features.
@@ 4 Non-theoretical Approaches of TDA in NLP
There are several ways to categorize the applied/non-theoretical TDA applications in NLP. These applications can be categorized by task, learning type, modality, TDA technique, and numerical representation. We observe that categorizing these TDA applications by task and numerical representation is more meaningful than the other categories since those categories are binary and not very descriptive of the landscape.
Out of these dimensions, the numerical representation showcases the bottleneck for extracting useful TDA features. See Figure 4 for an illustration of the pipeline for extracting TDA features from numerically represented texts. Therefore, while we focus on both task and numerical representation, our main taxonomy for the non-theoretical applications is centered on how TDA features are extracted from different forms of numerical representations. Figure 5 illustrates the taxonomy of non-theoretical applications of TDA in NLP tasks. See Table 5 in the Appendix for the non-theoretical approaches.
4.1 Tasks, Learning Types, Modality, and TDA techniques
@@
    Classification: The most popular application is deepfake text detection Løvlie (2023); Tulchinskii et al. (2024); Kushnareva et al. (2024, 2021); Uchendu et al. (2024); Wei et al. (2025).
@@
    Clustering and Topic Modeling: The most popular application is document clustering and topic modeling Holmes (2020); Guan et al. (2016).
@@
    Sentiment and Semantic Analysis: The most popular applications are linguistic/grammatical acceptability Cherniavskii et al. (2022); Jain et al. (2024), word sense induction and disambiguation Rawson et al. (2022); Temčinas (2018), and polysemy word classification Jakubowski et al. (2020); Shehu (2024).
@@
    Structure and Visualization: The most popular is using Mapper to visualize model hidden weights Garcia (2022); Rathore et al. (2023).
@@
    Health, Social, and Scholarly Analysis: Since this is not a popular application for TDA, the most interesting applications are - prediction of epidemics Petri and Leitao (2020) and categorization of lonely people Effah (2017).
@@
    Speech and Music Processing: The most popular applications are studying vocalizations Bonafos et al. (2023, 2024) and music classification Bergomi (2015); Sassone et al. (2022).
@@
    Model Interpretation and Analysis: The most popular applications are model probing to reveal behavior in hidden weights Kostenok et al. (2023); Gourgoulias et al. (2024).
@@
Finally, we can also categorize non-theoretical TDA applications in NLP by learning types, Modality, and TDA techniques. Learning types has supervised Elyasi and Moghadam (2019); Lavery et al. (2024), and unsupervised Spannaus et al. (2024); Bonafos et al. (2024); Modality has text Triki (2021); Kostenok et al. (2023), and Speech Sassone et al. (2022); and TDA techniques, has Persistent Homology Torres-Tramón et al. (2015); Cherniavskii et al. (2022), and Mapper Holmes (2020); Elyasi and Moghadam (2019).

### v4

@@ 4.1.1 Semantic Topological Space
A semantic topological space is a conceptual framework used to represent and analyze the relationships between the meanings (semantics) of words, phrases, or other linguistic units in a topological or shape structure. This representation involves mapping these units into a mathematical space where the distance or structure between them reflects semantic similarity or other relationships (i.e., Euclidean space $\to$ Topological space).
Karlgren et al. (2014) visualizes the topological semantic space of text using Mapper which identifies the topical density of the space. To capture topological properties, they train two semantic spaces in a specific topical domain (Karlgren et al., 2014). One space was trained only on articles of similar topics, and the other on introductory paragraphs of those same articles. Findings reveal that clusters of main concepts remained close for the space trained only on articles of similar topics. For the other topological space, the main concepts were randomly distributed (Karlgren et al., 2014). This suggests that richer and denser data can be used to capture the semantic topological space better than sparser data.
In addition, Cavaliere et al. (2017) extracts main concepts from the texts by probing the context-aware semantic topological space built with simplicial complexes. Finally, Wagner et al. (2012) uses TF-IDF to numerically represent the top 10-50 words in a corpus and build a topological space that analyzes the structure of similarities within several documents. This topological space is built using discrete Morse theory and Persistent Homology to find meaningful topological patterns (Wagner et al., 2012). However, as of 2012, they found that their technique was unsuccessful due to the computational costs, which is a testament to how the NLP field has improved such that we now have more tractable solutions, such as dimensionality reduction algorithms (McInnes et al., 2018), and TDA packages (Ripser (Bauer, 2021), Sklearn-TDA (Saul & Tralie, 2019), PHAT (Bauer et al., 2017), pytorch-topological [^4] https://github.com/aidos-lab/pytorch-topological), as well as compute resources to efficiently construct topological spaces from large or complex data.
@@ 4.1.2 Syntactic Topological Space
Therefore, Port et al. (2018) analyzes how syntactic parameters are distributed over different language families, including Indo-European, Niger-Congo, Austronesian, and Afro-Asiatic families. For instance, features in $\beta_{0}$ capture the subdivision into historical, and features in $\beta_{1}$ capture syntactic differences between branches of families of languages, as well as the syntactic influences between them (Port et al., 2018). They investigate the syntactic topological structures of language families, specifically Indo-European, Niger-Congo, Austronesian, and Afro-Asiatic families. Port et al. (2018) show that the three persistent connected components ($\beta_{0}$) in the Niger-Congo family represents its three subfamilies - Mande, Atlantic-Congo, and Kordofania. The syntactic topological structures of these languages also reveal the historical linguistic phenomena that the Hellenic branch played a role in the historical development of the Indo-European languages (Port et al., 2018).
Similarly, Port et al. (2022) probes the interpretability of the syntactic topological space. This is done by incorporating more explanation of linguistic structures by introducing so-called homoplasy phenomena to explain persistent loops further. Homoplasy phenomena in syntax are observed when dissimilar languages exhibit syntactic similarities (Port et al., 2022). Findings reveal that the Indo-European family languages - Czech, Lithuanian, Middle Dutch, and Swiss German have the same homoplasy phenomena (Port et al., 2022). Using this phenomenon to explain the appearance of persistent loops only when the 4 languages are present makes sense. This is because Middle Dutch and Swiss German are similar, however, Czech and Lithuanian are so different from them, making the homoplasy phenomenon the most reasonable explanation (Port et al., 2022).
Insight (Syntactic Space). The syntactic topological space captures the syntactic structure of language (i.e., grammar, etc.) from a topological lens. Using this framework, researchers confirm linguistic phenomena in language families and subfamilies by exploring the syntactic relationship between languages. Thus, a novel application of this framework could include the discovery of new linguistic phenomena within syntactic structures.
@@ 4.2 Topology of Topic Evolution
Thus, Sami & Farrahi (2017) utilizes TDA to visualize the relationship between words in a text block, words in a corpus, and text blocks in a corpus. Text blocks represent a chapter/section in a book, a document in a media corpus, and a webpage in a web corpus (Sami & Farrahi, 2017). They visualize both local context (i.e., each text block in a set of sentences), and global context (i.e., occurrence of extracted words in the corpus) features. These features are extracted by using the circular topology to represent words. Then, the peripheral nature of the text block and corpus can be visualized using these features. With the Local context features, dimension reduction is achieved by stemming the prefixes and suffixes of words. While, for the Global context features, word movement is captured, which analyzes topic evolution. Finally, findings reveal that using the circular topology in 2D space, core words from the corpus stay close to the center, and the explanatory words remain close to the circle’s periphery.
@@ 4.3 Topological “Shape” of Words
Draganov & Skiena (2024) captures the “shape” of words for several languages by comparing the phylogenies or evolutionary history of language in the Indo-European language family. Initially, numerically representing the texts with FastText, they use persistent homology to construct language phylogenetic trees for over 81 Indo-European languages. Experiments reveal that: (1) the shape of the word embedding of a language carries historical and structural information, similar to Port et al. (2018; 2022)’s findings; and (2) TDA methods can successfully capture aspects of the shape of language embeddings (Draganov & Skiena, 2024). Next, Fitz (2022) introduces a novel terminology - word manifold, which is a simplicial complex, whose topological space captures grammatical structure expressed by the corpus. This is done by implementing a technique for generating topological structure directly from strings of words. Experiments reveal that the homotopy type of the word manifold is also influenced by linguistic structure (Fitz, 2022). Similarly, Fitz et al. (2024) measures the topological complexity (known as perforation) of the hidden representation of Transformer-based models to understand their topological shapes.
Similar to Port et al. (2018; 2022) and Draganov & Skiena (2024), Dong (2024) extracts the topological shapes of languages. Specifically South American languages - the Nuclear-Macro-Jê (NMJ) and Quechuan families using TDA. By using techniques like multiple correspondence analysis (MCA) for dimension reduction of the categorical-valued dataset and persistent homology, Dong (2024) visualizes each language in the selected families as a point cloud. This forms the topological shape of the South American languages, such that languages close together are more similar. By comparing the topological shapes of the languages, it is observed that there are major distinctions between the Jê-proper and the non-Jê-proper languages, as well as the northern and southern Quechuan languages (Dong, 2024). Finally, Bouazzaoui et al. (2021) explores the topological similarity of the shapes of two writing systems - Tifinagh and Phoenician scripts.
@@
Name Task Problem Numerical Representation Learning Type Modality TDA
SIFT (Zhu, 2013) cl child vs. adolescent writing detection TF-IDF Supervised Text PH
(Løvlie, 2023) cl deepfake text detection TF-IDF, GloVe Supervised Text PH
(Huang, 2022) cl election speech feature extraction TF-IDF Supervised Text PH
(Doshi & Zadrozny, 2018; Shin, 2019) cl movie genre cl TF-IDF Supervised Text PH
(Sovdat, 2016) cl distinguishing between languages TF-IDF Supervised Text PH
(Savle et al., 2019) S & SA legal document entailment TF-IDF Supervised Text PH
(Michel et al., 2017) S & SA clustering and sentiment analysis TF-IDF, GloVe Supervised Text PH
DoCollapse (Guan et al., 2016) C & TM keyphrase extraction TF-IDF Unsupervised Text PH
TOPOL (Torres-Tramón et al., 2015) C & TM Twitter topic detection TF-IDF Supervised Text PH
(Kumar & Sarkar, 2022) C & TM text summarization TF-IDF Unsupervised Text PH
(van Veen, 2020) cl Propaganda tweet cl TF-IDF Unsupervised Text M
(Elyasi & Moghadam, 2019) cl classify Persian poems TF-IDF Supervised Text PH & M
(Effah, 2017) cl age group categorization of lonely people TF-IDF Supervised Text PH & M
(Maadarani & Hajra, 2020) cl nursery rhyme classification from different continents - Australia, Asia, Africa, Europe, and North America TF-IDF Supervised Text PH
(Haghighatkhah et al., 2022) S & V building document structure Pre-trained (Word2Vec) Unsupervised Text PH
BERT+TDA (Wu et al., 2022) S & SA contradiction detection Pre-trained (Word2Vec) Supervised Text PH
(Yessenbayev & Kozhirbayev, 2022) MI & A speaker recognition & text processing Pre-trained (Word2Vec) Unsupervised Speech PH
(Yessenbayev & Kozhirbayev, 2024) MI & A speaker recognition & text processing Pre-trained (Word2Vec) Unsupervised Speech PH
(Cornell, 2020) S & SA building a topological search engine Pre-trained (Word2Vec) Unsupervised Text M
(Holmes, 2020) C & TM document clustering and topic modeling tasks Pre-trained (Word2Vec) Supervised Text M
(Rawson et al., 2022) S & SA word sense induction and disambiguation Pre-trained (Word2Vec) Unsupervised Text PH
(Temčinas, 2018) S & SA word sense induction and disambiguation Pre-trained (Word2Vec, GloVe) Unsupervised Text PH
(Feng et al., 2024) MI & A Geometry of textual data augmentation Pre-trained (Word2Vec) Supervised Text PH
(Tymochko et al., 2021) cl fraudulent paper detection Pre-trained (Word2Vec, GloVe, ElMo) Supervised Text PH
(Petri & Leitao, 2020) H,S,& SA disease epidemic prediction Pre-trained (Word2Vec) Supervised Text PH
(Tymochko et al., 2020) cl finding topological loops in logical statements Pre-trained (Word2Vec) Unsupervised Text PH
(Wright & Zheng, 2020) C & TM distinguish subsets in data Pre-trained (Word2Vec) Unsupervised Text PH
(Paluzo Hidalgo et al., 2019) S & SA measuring the distance between the literary style of Spanish poets Pre-trained (Word2Vec) Supervised Text PH
(Bailey & Heiligman, 2025) S & SA detecting narrative shifts Pre-trained (Word2Vec) Unsupervised Text PH
(Gholizadeh et al., 2018) cl extract the topological signatures of novelists Pre-trained (GloVe) Supervised Text PH
TIES (Gholizadeh et al., 2020) S & SA document categorization & sentiment analysis Pre-trained (GloVe) Unsupervised Text PH
(Spannaus et al., 2024) MI & A phenotype prediction and news group categorization Pre-trained (GloVe) Unsupervised Text M
(Zadrozny, 2021b) S & SA finding topological loops in logical statements Pre-trained (GloVe) Unsupervised Text PH
(Byers, 2021) H,S,& SA social anxiety detection Pre-trained (GloVe) Supervised Text PH
(Haim Meirom & Bobrowski, 2022) MI & A compare cross-lingual sentence representations Pre-trained (GloVe) Unsupervised Text PH
(Zadrozny, 2021a) MI & A investigates the manifestations of intelligence and understanding in neural networks Pre-trained (GloVe) Supervised Text PH
(Deng & Duzhin, 2022) cl fake news detection Pre-trained (GloVe, BERT) Supervised Text PH
(Novak, 2019) H,S,& SA analyzing scholarly network Pre-trained (GloVe, BERT) Unsupervised Text PH
(Jakubowski et al., 2020) S & SA (1) polysemy word classification, and (2) word sense induction & disambiguation Pre-trained (FastText) Unsupervised Text PH
(Triki, 2021; Shehu, 2024) S & SA polysemy word cl Pre-trained (FastText) Supervised Text PH
PHD (Tulchinskii et al., 2024) cl deepfake text detection Pre-trained (Transformers - CLS) Supervised Text PH
Short-PHD (Wei et al., 2025) cl deepfake text detection (for short-text) Pre-trained (Transformers - CLS) Supervised Text PH
(Guilinger et al., 2025) cl deepfake text detection (for academic abstracts) Pre-trained (Transformers - CLS) Supervised Text PH
(Kushnareva et al., 2024) cl deepfake text detection Pre-trained (Transformers - CLS) Supervised Text PH
(Gourgoulias et al., 2024) MI & A class separability estimation Pre-trained (Transformers - CLS) Unsupervised Text PH
TopoBERT (Rathore et al., 2023) S & V and MI & A visually analyzing the fine-tuning process of a Transformer-based model Pre-trained (Transformers - CLS) Unsupervised Text M
(Lavery et al., 2024) cl fake news detection Pre-trained (Transformers - CLS) Supervised Text PH
(Das et al., 2021) cl classification of public speaking ratings from TED talks Pre-trained (Transformers - CLS) Supervised Text PH
(Byrne et al., 2022) C & TM topic modeling Pre-trained (Transformers - CLS) Unsupervised Text M
TOPFORMER (Uchendu et al., 2024) cl deepfake text detection Pre-trained (Transformers - Hidden) Supervised Text PH
(Garcia, 2022) S & SA polysemy word cl Pre-trained (Transformers - Hidden) Supervised Text M
(Bensalem et al., 2025) S & SA polysemy word cl Pre-trained (Transformers - Hidden) Unsupervised Text PH
(Alexander & Wang, 2023) H,S,& SA and S & V Hate speech, Misinformation & Psychiatric disorder cl Pre-trained (Transformers - Hidden) Supervised Text M
Persistent BERT Compression and Explainability (PBCE) (Balderas et al., 2025) MI & A Model compression Pre-trained (Transformers - Hidden) Unsupervised Text PH
Persistent Similarity (Gardinazzi et al., 2024) MI & A Probing layers in LLMs Pre-trained (Transformers - Hidden) Unsupervised Text PH
(Sun & Nelson, 2023) MI & A Correlation between sentence vectors Pre-trained (Word2Vec, Transformers - Hidden) Unsupervised Text PH
Persistence Scoring Function (Chauhan & Kaul, 2022) MI & A captures the homology of the high-dimensional hidden representations Pre-trained (Transformers - Hidden) Unsupervised Text PH
Topological Densification (García-Castellanos et al., 2024) MI & A zero-shot model stitching Pre-trained (Transformers - Hidden) Unsupervised Text PH
(Fitz, 2023) MI & A topology of fairness Pre-trained (Transformers - Hidden) Unsupervised Text M
(Fay et al., 2025) MI & A adversarial vs. non-adversarial text representation in LLMs Pre-trained (Transformers - Hidden) Unsupervised Text PH
(Ruppik et al., 2024) C & TM dialogue term extraction Pre-trained (Transformers - Hidden) Supervised Text PH
(Kudriashov et al., 2024) MI & A polypersonality Pre-trained (Transformers - Hidden) Unsupervised Text PH
(Bazarova et al., 2025) cl LLM Hallucination detection Pre-trained (Transformers - Attention) Supervised Text PH
(Kushnareva et al., 2021) cl deepfake text detection Pre-trained (Transformers - Attention) Supervised Text PH
(Cherniavskii et al., 2022; Proskurina et al., 2023; Jain et al., 2024) S & SA grammatical acceptability judgment Pre-trained (Transformers - Attention) Supervised Text PH
(Kostenok et al., 2023) MI & A Uncertainty estimation of model predictions Pre-trained (Transformers - Attention) Supervised Text PH
(Perez & Reinauer, 2022) cl spam detection, grammatical acceptability judgment, and movie sentiment analysis Pre-trained (Transformers - Attention) Supervised Text PH
(Sakurai et al., 2025) cl Authorship attribution of Japanese texts Pre-trained (Transformers - Attention) Supervised Text PH
(Pollano et al., 2024) cl out-of-distribution detection Pre-trained (Transformers - Attention, CLS) Supervised Text PH
(Proskura & Zaytsev, 2024) MI & A estimation of weights for ensembles of classification models Pre-trained (Transformers - Attention, CLS) Supervised Text PH
(Arun et al., 2025) S & V and S & SA controversial vs. non-controversial political discourse detection Pre-trained (Transformers - CLS) Supervised Text PH
(Snopov & Golubinskiy, 2024) cl vulnerability detection in code Pre-trained (Transformers - Attention) Supervised Text PH
TopoHuBERT (Tulchinskii et al., 2023) S & MP speaker recognition Pre-trained (Transformers - Attention) Supervised Speech PH
(Vukovic et al., 2022) S & SA dialogue term extraction Pre-trained (Transformers - Attention) Supervised Text PH
(Yurchuk & Gurnik, 2023) S & SA Ukrainian tongue twisters cl Symbolic Representations Supervised Text PH
(Kovaliuk et al., 2024) S & MP Ukrainian tongue twisters cl Symbolic Representations Supervised Speech PH
(Bonafos et al., 2023) S & MP human vowel cl Multi-Modal Supervised Speech PH
(Bonafos et al., 2024) S & MP infant vocalization cl Multi-Modal Unsupervised Speech PH
(Gonzalez-Diaz et al., 2019; Paluzo-Hidalgo et al., 2022) S & MP emotion recognition Multi-Modal Supervised Speech PH
(Tlachac et al., 2020) S & MP depression detection Multi-Modal Supervised Speech PH
(Zhu et al., 2024) S & MP consonants recognition Multi-Modal Supervised Speech PH
(Sassone et al., 2022) S & MP music genre cl Multi-Modal Supervised Speech PH
(Vu et al., 2025) S & MP multi-modal adversarial robustness assessment Multi-Modal Supervised Text PH
(Bergomi, 2015) S & MP music cl Multi-Modal Supervised Speech PH
Table 3: Non-theoretical applications of TDA in NLP. For the TDA techniques - PH: Persistent Homology and M: Mapper. Task categories - cl: classification, C & TM: clustering & topic modeling, S & SA: sentiment & semantic analysis, S & V: structure & visualization, H,S,& SA: health, social, & scholarly analysis, S & MP: speech & music processing, MI & A: model interpretation & analysis.
@@
Figure 6: Illustration of the Non-Theoretical approach of using TDA as feature extractor in NLP with three steps: (1)-extracting numerical representations, (2)-reformatting for TDA’s inputs, and (3)-extracting TDA features.
@@ 5 Non-theoretical Approaches of TDA in NLP
There are several ways to categorize the applied/non-theoretical TDA applications in NLP. These applications can be categorized by task, learning type, modality, TDA technique, and numerical representation. We observe that categorizing these TDA applications by task and numerical representation is more meaningful than the other categories since those categories are binary and not very descriptive of the landscape. Out of these dimensions, the numerical representation showcases the bottleneck for extracting useful TDA features. See Figure 6 for an illustration of the pipeline for extracting TDA features from numerically represented texts. Therefore, while we focus on both task and numerical representation, our main taxonomy for the non-theoretical applications is centered on how TDA features are extracted from different forms of numerical representations. Figure 4 illustrates the taxonomy of non-theoretical applications of TDA in NLP tasks. See Table 3 for the list of non-theoretical approaches.
Figure 7: Number of publications in the sub-categories for the non-theoretical Taxonomies - Numerical Representations, Tasks, Modalities, and TDA techniques categories.
5.1 Tasks
@@
    Classification: The most popular application is deepfake text detection (Løvlie, 2023; Tulchinskii et al., 2024; Kushnareva et al., 2024; 2021; Uchendu et al., 2024; Wei et al., 2025).
@@
    Clustering and Topic Modeling: The most popular application is document clustering and topic modeling (Holmes, 2020; Guan et al., 2016).
@@
    Sentiment and Semantic Analysis: The most popular applications are linguistic/grammatical acceptability (Cherniavskii et al., 2022; Jain et al., 2024), word sense induction and disambiguation (Rawson et al., 2022; Temčinas, 2018), and polysemy word classification (Jakubowski et al., 2020; Shehu, 2024).
@@
    Structure and Visualization: The most popular is using Mapper to visualize model hidden weights (Garcia, 2022; Rathore et al., 2023).
@@
    Health, Social, and Scholarly Analysis: Since this is not a popular application for TDA, the most interesting applications are - prediction of epidemics (Petri & Leitao, 2020) and categorization of lonely people (Effah, 2017).
@@
    Speech and Music Processing: The most popular applications are studying vocalizations (Bonafos et al., 2023; 2024) and music classification (Bergomi, 2015; Sassone et al., 2022).
@@
    Model Interpretation and Analysis: The most popular applications are model probing to reveal behavior in hidden weights (Kostenok et al., 2023; Gourgoulias et al., 2024).
@@
5.2 Other Taxonomies: Learning Types, Modality, and TDA techniques
Finally, we can also categorize non-theoretical TDA applications in NLP by learning types, Modality, and TDA techniques. Learning types has supervised (Elyasi & Moghadam, 2019; Lavery et al., 2024), and unsupervised (Spannaus et al., 2024; Bonafos et al., 2024); Modality has text (Triki, 2021; Kostenok et al., 2023), and Speech (Sassone et al., 2022); and TDA techniques, has Persistent Homology (Torres-Tramón et al., 2015; Cherniavskii et al., 2022), and Mapper (Holmes, 2020; Elyasi & Moghadam, 2019).
Lastly, see Figure 7 for the number of publications for each sub-category of the four taxonomies - Numerical Representation, Task, Modality, and TDA techniques.